OpenStudy (anonymous):

The function f is defined by f(x)=2x^2 -1 . Find f(4x). I'm really having trouble understanding what this question wants. I got 32x^2 -1 but I don't think i did it right.

OpenStudy (alyssa_xo):

replace x with (4x) \(f(4x)=2(4x)^2-1=2*16x^2-1=32x^2-1\) you did it right

OpenStudy (anonymous):

okay. Thank you. I just wasn't sure if i replace it with 4 or 4x. Thank you! If I need more help can I tag you?

OpenStudy (alyssa_xo):

sure

OpenStudy (anonymous):

Thank you. Every time I post questions no one helps, this site has failed me so many times.

OpenStudy (anonymous):

I'll close the question now.

OpenStudy (alyssa_xo):

it depends on a lot of stuff, name time of day, and what day it is. on weekends/fridays nobody's really on

OpenStudy (alyssa_xo):

alright

OpenStudy (alyssa_xo):

namely*